# CounterQuill: Investigating the Potential of Human-AI Collaboration in Online Counterspeech Writing

### v1

@@ 6.3.2. Developing Empathetic Reflection through Reflective Questions
Additionally, as mentioned by some participants (6 — P2, P7 - P9, P16, P20), the question-answering process guided them to engage in empathy-based counterspeech by considering the harmful impact and feelings of the targeted group. P2 elaborated on this shift in perspective:
@@
In addition to the questions-answering, most participants (16 — P1-P7, P9-P12, P15, P17, P19, P20) also described benefiting from the system’s feedback on their response reflection and suggested counterspeech strategies (§3.2 System Design - Brainstorming). These features helped reduce their anxiety and mental load, particularly when they were unsure how to begin writing counterspeech. P5 specifically noted that the system’s suggestions alleviated the stress of starting counterspeech from scratch.
@@
Participant 5: After I got the feedback, there were things I wanted to add to my comments or input. It was like, ’Oh, I could add that.’ The feedback gave me a good starting point for writing counterspeech. If you asked me to write a counterspeech from scratch, I would start from a place of uncertainty, and it might not be very straightforward. But the feedback gave me a clear target to move on.
@@
6.4. Qualitative Findings of Writing with CounterQuill
6.4.1. Empowering Users through CounterQuill’s Writing Session: From Brainstorming to Confident Counterspeech
Most participants commented on the intuitive design and user-friendliness of CounterQuill’s writing interface. The friendly layout and convenient features allowed participants to focus on composing their counterspeech without being hindered by complex technology. P2 summarized his experience:
[…]
Participant 2: The writing interface is very user-friendly and easy to use. It allowed me to concentrate on writing my response without getting distracted by complicated features.
@@ 6.4. Qualitative Findings of Writing with CounterQuill
6.4.2. Co-writing with CounterQuill: Selective Rewriting, Improved Confidence, and Increased Ownership
@@
Participant 1: … I think that’s a lot better than ChatGPT because if you want to do selective edits with ChatGPT, you have to tell it specifically that you don’t want to change certain parts, or you would have to take those parts out and then only insert the part that you want to edit. But with this tool, you can just do the selective rewriting in one place, which is an improvement.
In addition to selective rewriting, participants mentioned the features of CounterQuill’s writing assistant. Among these, many participants (12 — P2 - P6, P10, P11, P12, P17 - P20) reported that grammar checking was the first feature they would use. All participants (20) had thoughts on the empathetic rewriting feature: the empathetic rephrasing suggestions helped participants better express their ideas (17 — P1 - P12, P15 - P19), making their responses more relatable and humanizing (17 — P1 - P8, P10 - P13, P15 - P18, P20). The empathetic rephrasing suggestions helped participants better articulate their thoughts (11 — P2, P6 - P13, P19, P20), enhanced the impact of their counterspeech (15 — P1, P2, P3 - P9, P11 - P14, P18, P20) and give them emotional support (7 — P2, P3, P5, P7, P12, P14, P18). For example, P2 felt more connected and less confrontational in writing counterspeech and shared how the system improved the original counterspeech:
@@
[…]
Participant 14: I tried many prompts (with ChatGPT), but even if it gave me 10 different counterspeech, I wouldn’t use any of them … because they are too AI-like.
In contrast, some participants expressed a preference for continuing to use CounterQuill’s writing features in future real-world online interactions. They particularly appreciated how CounterQuill integrated prompts like tone adjustment and rewriting into convenient buttons, which made the counterspeech writing process more efficient than using other chatbots (“I am likely to use CounterQuill in the future because its system features, like changing tone, help me write counterspeech more quickly than some chatbots” - P2).
@@ 7. DISCUSSION
7.1. Design Implications for AI-Mediated Counterspeech Tools
Designing AI tools that cater to a wide range of individuals is essential, particularly in social media environments where diverse attitudes towards counterspeech are prevalent. Prior work has highlighted the varying motivations, barriers, and skills among different groups when countering online hate speech (Mun et al., 2024; Ping et al., 2024; Zhu and Bhat, 2021). This diversity prompts the question: how can we design customized AI tools that effectively function in social media environments, where users with diverse attitudes engage in counterspeech? How can these tools influence the thought patterns and behaviors of such varied groups?
Our study provides insights to help address these questions in the context of improving ways to engage in effective counterspeech. Our findings revealed varied attitudes from participants when faced with the challenge of writing counterspeech. For instance, P1 expressed a strong moral obligation to write counterspeech (‘well I think about it this way, if someone is saying something nasty and you are silent, it’s like letting the bad stuff win. I just feel like engaging in counterspeech is the right thing to do.’ ), while P4 hesitated to participate due to fear of potential backlash or lack of confidence in writing abilities. For instance, when asked about how confident they are in writing counterspeech before using CounterQuill, P4 expressed concerns— ‘I am always worried about what might happen if I speak back. What if people attack me for it? Also, I’m not sure if I can speak my thoughts well enough.’ Such varying attitudes and reactions reinforce us to design AI-assisted tools that can provide tailored suggestions to users, aligning with their unique comfort levels and engagement preferences in countering hate speech. Future research could explore the potential of embedding tools like CounterQuill directly within social media platforms (e.g., Facebook, Instagram, Reddit, Twitter) as extensions for browsers. Such an approach could make counterspeech tools more accessible and convenient for users, reducing the time spent navigating web interfaces and allowing for real-time, context-sensitive suggestions. By focusing on the balance between time and outcomes, integrating features like tone adjustment and rewriting prompts may provide users with immediate, relevant guidance based on their interaction history, such as posts and comments.
Moreover, as P5 stated - “The highlighting exercise made it easy to get the main points [in the hate speech]”, users can benefit from AI offering more personalized explanations. For instance, as our findings show, these explanations can focus on specific elements like the identity of the targeted individual/ group or the dehumanizing actions to help users comprehend and respond to hate speech effectively.
In addition, participants were inconsistent in their understanding of counterspeech. For example, participants (P4, P18) recognized their limitations in responding to hate speech but were not aware of specific terminology (‘I did respond to some hate [speech] before, but I never [knew] the term (counterspeech)’), while others were unaware of the terminology altogether (P5, P7). This disparity in user perceptions extended to their goals and motivations in engaging in counterspeech. Therefore, AI-assisted counterspeech research should address users’ varied needs and backgrounds engaging with these tools, particularly LLM-powered systems. To enhance counterspeech efforts, these tools should prioritize adaptability and personalization, striking a balance between efficiency and effectiveness. AI-mediated tools should aim to provide tailored guidance based on users’ personal experiences and specific goals in engaging with counterspeech.
7.2. AI-Mediated Communication for Counterspeech Ownership & Engagement
With the advancement of AI technology, AI-mediated communication (AIMC) offers new possibilities for assisting individuals in overcoming these barriers and writing counterspeech (Mun et al., 2024; Zhu and Bhat, 2021; Ping et al., 2024). However, the introduction of AIMC has also raised concerns, particularly regarding participants’ sense of ownership of the AI-generated content, which is related to their engagement with counterspeech (e.g., whether or not to post counterspeech on social media).
In this section, we discuss potential outcomes that may arise from AI-mediated counterspeech, such as participants’ sense of ownership over the co-written counterspeech and their level of engagement.
Figure 12. The relationship between user effort, sense of ownership, and willingness to post counterspeech on social media: more cognitive effort invested in the AI-mediated system leads to a greater sense of ownership and increased likelihood of posting the co-created content online. The blue background area is indicated: past research has demonstrated the correlation between cognitive effort and ownership of human AI co-created content.
7.2.1. How The Sense of Ownership Manifested through Counterquill Usage
Our findings discovered that when participants engage in collaborative writing with CounterQuill, they develop a sense of ownership over the counterspeech. This occurs when participants incorporate personal experiences and values into their writing process and make revisions based on CounterQuill’s suggestions. Our findings echo previous HCI research, (Figure 12), showing that the more meaningful cognitive effort people invest in collaborating with AIMC systems, the stronger their sense of ownership becomes (Louie et al., 2021, 2020a; Pierce et al., 2003; Mieczkowski, 2022).
Through our collaboration with CounterQuill, we found that participants were surprised by how the system co-authored counterspeech with them. Many referred to CounterQuill as a writing “partner” or “friend” rather than simply an AI or model. We believe the following features of the system contributed to the increased sense of ownership. First, participants could maintain control by selecting AI-generated text suggestions or querying the system for tone modifications using the “AI Writing Assistant” interface. Second, CounterQuill preserved their efforts, unique voices, and tones. The system’s generated results maintained the context and semantics of the user’s original content, aligning with their intentions and personal experiences. Last, participants in the human-AI collaboration condition could actively insert and delete personal content during the writing process. They engaged in thoughtful deliberation while writing and reported the feeling of initiative and control over the writing process, leading to a stronger sense of ownership.
In addition, we found that the collaborative nature of the co-writing process promoted a sense of shared goals, echoing Caspi et al.’s findings. (Caspi and Blau, 2011). Participants perceived the final counterspeech as “ours” rather than merely “mine,” further enhancing their sense of ownership. Specifically, many participants in our study believed that the final response was completed by “CounterQuill and me together.” This co-ownership emerged when participants invested cognitive effort, energy, and even personal values, and the AI system provided suggestions and guidance that aligned with the participants’ experiences and personal values.
7.2.2. Comparing ChatGPT and CounterQuill : Ownership and Engagement
However, the sense of ownership over information is not without complexity. People tend to believe that the results of co-creation with AI belong to “us”, while simultaneously acknowledging that AI-generated text is trained from the language of potentially billions of others. This tension, as suggested by Mieczkowski (2022), may lead to a slight defensiveness in some individuals due to the perceived threat to their sense of ownership (Mieczkowski, 2022).
Through our study, we discovered that participants’ sense of ownership over counterspeech was diminished during the baseline task. Most participants pointed out that “CounterQuill is different from ChatGPT,” because people typically invest less time and interaction with conversational AI tools (i.e., ChatGPT), resulting in a lower sense of ownership. They often describe these tools as “assistants” rather than “collaborators”. This finding echoes Louie et al. and Mieczkowski’s study (Figure 12), where people expressed a lower sense of ownership due to their limited investment in the task (Mieczkowski, 2022; Louie et al., 2020a), even describing themselves as “lazy” (Louie et al., 2020a, b).
Furthermore, our study found that during interactions with ChatGPT, participants felt “passive” (Dummer et al., 2009; Synofzik et al., 2008) because they could neither predict the output nor change it after insertion. Consequently, participants opted for “lazy prompts” when conversing with ChatGPT. However, even when individual participants attempted to write elaborate prompts, they experienced resistance (“I tried many prompts, but even if it gave me 10 different counterspeech, I wouldn’t use any of them, because they are too AI-like” - P14). These findings are consistent with previous studies (Louie et al., 2020a; Kim and Song, 2023), indicating that human interactions with conversational AI, such as ChatGPT, lack the back-and-forth exchange and the investment of personal effort. Consequently, people generally express a weak sense of ownership over the results generated by these AI “assistants” and refrain from using them.
7.2.3. The Relationship Between Ownership and Willingness to Share
Our study reveals that participants’ perception of ownership and the amount of time (perceived effort) they invest in CounterQuill influence their end goal of whether they will post the counterspeech online.
Our quantitative analysis (§6.1) showed that collaborating with ChatGPT required less cognitive effort than CounterQuill. However, in our customized questions, participants reported higher satisfaction with the final counterspeech generated in collaboration with CounterQuill, with an average user rating of 6.47, significantly higher than ChatGPT (4.00) by 62%. Moreover, participants were more inclined to publish the final generated content online when collaborating with CounterQuill, with an average user rating of 6.29, considerably higher than ChatGPT (3.53) by 78%. Although our user study did not directly measure perceived ownership of the generated content, we found that participants’ satisfaction and willingness to publish online, which can be seen as a proxy of perceived ownership, were higher in CounterQuill’s condition. This finding echoes previous research by Lee et al. (2015), who discovered that people are willing to share publicly for various motivations, for example, sharing their own ideas and speeches (“To express my actual self” and “To share my personal information with others”) (Lee et al., 2015).
Combining our qualitative analysis, our findings indicate that when working with AI, people tend to develop a strong sense of ownership over co-created goals due to the investment of time, energy, and even personal values required in the process of achieving these goals. Consequently, they are more willing to publish content that belongs to ”ours” on social media. However, when AI-generated content is produced quickly and with minimal effort, the sense of ownership loss becomes particularly evident. Furthermore, when users consider the potential perspectives of others - how the recipients of the information might view the sender’s use of AI-generated text - they are less likely to publish it on social media.
@@ 8. CONCLUSION
In this work, we introduced CounterQuill, an AI-mediated system designed to help users write effective counterspeech. CounterQuill guides users through a learning session to understand hate speech and counterspeech concepts, a brainstorming session to identify hateful elements and generate response ideas, and a co-writing session to draft and refine their counterspeech.
Our user study revealed that CounterQuill’s learning session effectively educated users about hate speech and empathy-based counterspeech strategies. The brainstorming session, with its AI guidance and feedback, improved users’ confidence and perceived ability to engage in counterspeech. Furthermore, the co-writing session, which provided AI writing assistance and personalization options, enhanced users’ self-efficacy and ownership of the counterspeech they created. Interestingly, we found that the more meaningful effort users invested in collaborating with CounterQuill, the stronger their sense of ownership over the co-created content, leading to a higher likelihood of posting the counterspeech on social media.
Limitations and future work. While our research provides insights into the potential of using AI-mediated system for countering online hate speech, further research is needed to explore these nuances of human-AI collaboration and the factors that influence users’ willingness to participate. For example, past research found that individuals who confronted those who were engaging in hate speech often became targets themselves (Costello et al., 2017; Hawdon et al., 2017). This victimization, in turn, can reduce one’s willingness to participate if the person sees hate speech in the future (Hawdon and Costello, 2022). One limitation of our user study is that participants’ prior experience and competence with AI and counterspeech may have influenced their performance and perceptions when using CounterQuill and ChatGPT, potentially impacting the study’s results. Additionally, our study was restricted to participants responding to a single hate post. Future research should include multiple hate posts to gain more comprehensive insights and better assess CounterQuill’s effectiveness across different contexts.
Future research should investigate if using AI-generated counterspeech reduces the likelihood of the user being targeted or provides the user greater confidence to generate counterspeech even if they are victimized. If human-AI collaboration can both improve users’ sense of ownership and reduce their apprehension about participating in counterspeech, we will be able to design more effective tools to combat online hate speech and foster healthier online communities.
Building on this, we propose developing CounterQuill as a Chrome or Edge extension that can be integrated directly into social media platforms like Reddit and Twitter. Our future work will include a long-term user study of the CounterQuill extension, where participants will use it in their daily lives to identify hate speech and collaborate with AI to compose counterspeech responses on real social media platforms. We also plan to conduct more granular comparisons to showcase the unique writing features of the CounterQuill extension with other AI tools.
Appendix A pre-screening survey

### v2

@@ 6.3.2. Stimulating Reflective Ideation Through Guided Questions
Additionally, as mentioned by some participants (6 — P2, P7 - P9, P16, P20), the question-answering process guided them to consider the harmful impact and feelings of the targeted group. P2 elaborated on this shift in perspective:
@@
In addition to the questions-answering, most participants (16 — P1-P7, P9-P12, P15, P17, P19, P20) described benefiting from the system’s feedback on their response reflection and suggested counterspeech strategies (§4.2.2). These features helped reduce their anxiety and mental load, particularly when they were unsure how to begin writing counterspeech. P5 specifically noted that the system’s suggestions alleviated the stress of starting counterspeech from scratch.
@@
Participant 5: After I got the feedback, there were things I wanted to add to my comments or input. It was like, ‘Oh, I could add that.’ The feedback gave me a good starting point for writing counterspeech. If you asked me to write a counterspeech from scratch, I would start from a place of uncertainty, and it might not be very straightforward. But the feedback gave me a clear target to move on.
@@
6.4. Insights from the Co-Writing Session
6.4.1. Transforming Brainstormed Ideas into Counterspeech
Most participants commented on the intuitive design and user-friendliness of CounterQuill’s writing interface. “The friendly layout and convenient features allowed participants to focus on composing their counterspeech” (P2), without being hindered by complex technology.
@@ 6.4.1. Transforming Brainstormed Ideas into Counterspeech
6.4.2. Achieving a Sense of Ownership Through Collaborative Revision
@@
In addition to selective rewriting, many participants (12 — P2 - P6, P10, P11, P12, P17 - P20) reported that grammar checking was the first feature they would use. All participants (20) had thoughts on the empathetic rewriting feature: the empathetic rephrasing suggestions helped participants better express their ideas (17 — P1 - P12, P15 - P19), making their responses more relatable and humanizing (17 — P1 - P8, P10 - P13, P15 - P18, P20). The empathetic rephrasing suggestions helped participants better articulate their thoughts (11 — P2, P6 - P13, P19, P20), enhanced the impact of their counterspeech (15 — P1, P2, P3 - P9, P11 - P14, P18, P20) and give them emotional support (7 — P2, P3, P5, P7, P12, P14, P18). For example, P2 felt more connected and less confrontational in writing counterspeech and shared how the system improved the original counterspeech:
@@
While some participants found the process of generating counterspeech with ChatGPT to require less effort (e.g., “ChatGPT can provide a quick response” - P14), they expressed a lack of ownership over the content generated by ChatGPT and were hesitant to post AI-generated counterspeech online (“I tried many prompts (with ChatGPT), but even if it gave me 10 different counterspeech, I wouldn’t use any of them … because they are too AI-like”). In contrast, participants (19 — P1-P19) preferred CounterQuill for future real-world online interactions. They particularly appreciated how CounterQuill incorporates tone adjustment while preserving their own voice and personal experiences (“I am likely to use CounterQuill in the future because its system features really keep my own word.” - P2).
@@ 7. DISCUSSION
Our findings provide insights for designing AI-assistant tools that empower users in online counterspeech writing. We discuss three key design implications in this context.
7.1. Encouraging User Ownership in AI-Assisted Writing
A major challenge in AI-mediated writing tools is maintaining a sense of personal ownership over AI-generated content (Louie et al., 2020; Mieczkowski, 2022). Although our study did not directly measure perceived ownership of the generated content, we found that participants’ satisfaction and willingness to post their written counterspeech online, which can be seen as a proxy of perceived ownership, were higher in CounterQuill’s condition. These findings echo prior research by Bhattacherjee et al. (2001), who found that higher levels of satisfaction create a sense of ownership of the results generated by the system, which in turn drives users to continue using the system (Bhattacherjee, 2001). When participants actively modified AI-generated content rather than simply accepting it, they reported a stronger sense of ownership over the counterspeech (§ 6.4.2). This was further reinforced when users incorporated personal experiences into their counterspeech and adjusted the AI’s suggestions to match their intended tone. Participants described the process as ”co-writing with the AI,” with some even referring to CounterQuill as a writing partner rather than just an automated tool.
Our work shows that integrating user-driven ideation and revision in AI-assisted counterspeech writing can increase users’ sense of ownership over their counterspeech. Prior research in HCI has shown that the more meaningful cognitive effort people invest in collaborating with AIMC systems, the stronger their sense of ownership towards the final output (Louie et al., 2021, 2020; Pierce et al., 2003; Mieczkowski, 2022). Our findings echo this: participants who invested effort in refining the AI’s suggestions—rather than passively accepting them—expressed a stronger satisfaction with their counterspeech. By designing AI assistance to prioritize user engagement in the writing process, AI-mediated communication (AIMC) tools like CounterQuill can empower users to take ownership of their responses rather than feeling as though they are merely transmitting AI-generated content.
Similarly, as highlighted in (§ 6.4.1), participants who had explicit control over CounterQuill’s customization settings were more likely to engage with and revise AI-generated suggestions. Research has shown that while AI-assisted writing can improve efficiency, excessive reliance on automation reduce user agency and engagement (Ping et al., 2024; Lee et al., 2022, 2024a). Our findings show that users value AI assistance that enhances their existing ideas rather than replacing them. To achieve this, AI writing assistants for counterspeech engagement should prioritize user-driven customization (Zhang et al., 2023), allowing users to selectively refine, rephrase, or adjust AI suggestions rather than accepting predefined outputs. Furthermore, embedding transparency mechanisms (e.g., visual indicators of AI contributions (Akhtar, 2024)) can help users retain control over their writing process, fostering trust and a greater sense of ownership (Mieczkowski, 2022; Draxler et al., 2024; Kim and Song, 2023).
7.2. Structuring AI Guidance to Reduce Cognitive Load
Effective counterspeech requires careful thought, yet users often experience cognitive overload when crafting responses to hate speech (Ping et al., 2024). CounterQuill’s step-by-step workflow—incorporating learning, brainstorming, and structured AI assistance—helps mitigate this challenge by breaking down complex tasks into manageable stages (§ 6.4). Echoing prior research (Ping et al., 2024), our findings suggest that AI writing assistants should avoid overwhelming users with unstructured suggestions; instead, they should provide guided interactions that progressively build toward the final output (Ping et al., 2024; Zhang et al., 2023).
Rather than providing fully generated responses, AI systems should support incremental co-creation, allowing users to shape and refine content instead of passively receiving suggestions (Caspi and Blau, 2011). Features like selective rewriting, interactive brainstorming, and context-aware feedback as demonstrated in CounterQuill help users integrate their own voice and perspective, increasing their willingness to engage in online discussions. By guiding users through the writing process, structured AI assistance can build confidence and encourage more deliberate, personalized counterspeech.
However, this structured guidance comes with a trade-off. CounterQuill’s multi-step workflow requires more time and effort than fully automated approaches like ChatGPT. While this level of engagement benefits users seeking to learn, brainstorm, and refine their responses (§ 6.4), it may not be suitable for experienced counterspeech writers who may want to prioritize speed to counteract and drown out online hate speech as quickly as possible before it spreads (Schieb and Preuss, 2016; ling Chung et al., 2023; Ping et al., 2024).
Our quantitative analysis (§6.1) showed that collaborating with ChatGPT required less cognitive effort and time than CounterQuill. However, in customized post-task questions, participants reported higher satisfaction with their final counterspeech when using CounterQuill, rating it 6.47 compared to ChatGPT’s 4.00—a 62% increase. Similarly, participants were 78% more likely to publish their counterspeech online when collaborating with CounterQuill (6.29 vs. 3.53). These findings suggest that while structured AI guidance demands more effort, it fosters deeper engagement and a stronger sense of ownership—key factors in promoting sustained participation in counterspeech efforts. Striking a balance between structured support and efficiency will be essential for designing AI systems that cater to both novice and experienced counter speakers.
7.3. Enhancing AI’s Role in Online Discourse Moderation
As platforms shift toward community-driven moderation (Brewer et al., 2020), AI-assisted counterspeech tools like CounterQuill can serve as more than just writing aids - they can actively shape how users engage with harmful content. Rather than replacing human content moderation with automated enforcement, AI can equip users with strategic communication skills to intervene in online discussions more effectively. This approach complements content moderation policies by fostering peer-led counter-messaging rather than relying solely on content removal.
Our findings indicate that structured AI guidance not only reduces cognitive load (§ 6.4.1) but also increases users’ confidence in their responses (§ 6.4.1), making them more willing to engage in public discourse. Participants appreciated AI-generated guidance that included contextual explanations alongside counterspeech suggestions, reinforcing their ability to engage in meaningful dialogue. Moreover, our results suggest that the greater the effort users invested in co-creating counterspeech, the more likely they were to share it publicly. This suggests that AI-mediated tools, when designed to prioritize user agency and strategic guidance, can encourage more active and constructive interventions in online discussions that may be fraught with hateful messages.
Finally, future AI-mediated moderation tools should focus not just on detecting and removing harmful content, but on educating and empowering users to respond constructively. By integrating adaptive AI guidance that responds to user expertise levels, these systems can help bridge the gap between passive content filtering and scalable, community-driven intervention strategies.
@@ 8. CONCLUSION
In this work, we introduced CounterQuill, an AI-mediated system designed to assist users in writing effective counterspeech. CounterQuill guides users through three key phases: a learning session to enhance understanding of hate speech and counterspeech, a brainstorming session to identify hateful elements and generate response strategies, and a co-writing session to refine counterspeech collaboratively with AI assistance.
Our user study demonstrated that CounterQuill effectively educates users about hate speech and empathy-driven counterspeech strategies. The brainstorming session, supported by AI feedback, increased users’ confidence in engaging with counterspeech, while the co-writing session enhanced their sense of self-efficacy and ownership. Notably, we found that the more effort users invested in collaborating with CounterQuill, the stronger their sense of ownership over the co-created content—leading to a higher likelihood of posting their responses on social media.
Limitations and future work. One limitation of our study is the potential influence of participants’ prior experience with AI and counterspeech on their engagement with CounterQuill. Additionally, our evaluation was limited to responses to a single hate speech example, which may not fully capture the system’s adaptability across different contexts. Future work should explore multiple hate speech scenarios to assess CounterQuill’s effectiveness more comprehensively.
Although participants indicated they would use CounterQuill to craft counterspeech, perceived effectiveness does not always translate into real-world adoption. However, the Theory of Planned Behavior (Ajzen, 1980, 1991) suggests that positive perceptions of a tool are strong predictors of future usage (Wan et al., 2014; Bosnjak et al., 2020). Given that users are unlikely to adopt a tool they perceive as ineffective, our findings indicate that CounterQuill holds promise for real-world impact.
For future research, we propose developing CounterQuill as a browser extension (e.g., for Chrome and Edge) and integrating it into social media platforms like Reddit and X. Additionally, we plan to conduct a longitudinal study to evaluate CounterQuill extension’s usability and effectiveness in real-world online discourse.
Appendix A Preliminary survey questions